OpenStudy (anonymous):

A circle is shown on the coordinate plane below. Circle that contains points 3, 0; 1, 2; negative 1, 0; 1, negative 2. What is the equation of this circle? (x – 1)2 + y2 = 2 (x + 1)2 + y2 = 4 (x + 1)2 + y2 = 2 (x – 1)2 + y2 = 4

OpenStudy (anonymous):

Please help, will give a medal to best answer

jimthompson5910 (jim_thompson5910):

Hint: Let A be the point (-1,0) and B be the point (3,0) Now find the midpoint of AB. This will give you your center.

OpenStudy (anonymous):

I have no idea how to find a midpoint, I'm not looking for an anwer as much as I'm looking for an explaination.

jimthompson5910 (jim_thompson5910):

example midpoint of (6,7) and (8,9) add up the x coordinates and divide by 2 x = (6+8)/2 = 14/2 = 7 add up the y coordinates and divide by 2 y = (7+9)/2 = 16/2 = 8 So the midpoint in this example is (7,8)

OpenStudy (amistre64):

you can try to plug inthe point values into the given equations to see which one fits the best.

OpenStudy (amistre64):

otherwise, it hard to parse the information that you have preseneted without going into alot of manual labor :)

OpenStudy (anonymous):

okay so we found the midpoint, what now?

jimthompson5910 (jim_thompson5910):

what did you get for the midpoint

OpenStudy (anonymous):

I didn't do it, can you just show me in depth?

OpenStudy (anonymous):

-2 /2 = -1?

jimthompson5910 (jim_thompson5910):

A is the point (-1,0) and B is the point (3,0) midpoint x = (-1+3)/2 x = 2/2 x = 1 y = (0+0)/2 y = 0/2 y = 0 midpoint of AB is (1,0)

OpenStudy (anonymous):

okay, so our midpoint is 1,0; now what?

jimthompson5910 (jim_thompson5910):

that's the center of the circle

jimthompson5910 (jim_thompson5910):

the distance from the center (1,0) to either A or B is what?

OpenStudy (anonymous):

so aslong as you find the midpoint, you find the center of the circle.

OpenStudy (anonymous):

where is A, and B

jimthompson5910 (jim_thompson5910):

A(-1,0) and B(3,0)

OpenStudy (anonymous):

the distance would be 2

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

so the radius is 2

OpenStudy (anonymous):

okay, so now we know our radius. Now what?

jimthompson5910 (jim_thompson5910):

(x – h)^2 + (y-k)^2 = r^2 (x – 1)^2 + (y-0)^2 = 2^2 (x – 1)^2 + y^2 = 4

jimthompson5910 (jim_thompson5910):

(h,k) is the center, r is the radius

OpenStudy (anonymous):

jim_thompson5910, you are very very smart I appreciate all your help. Thank you sooooooooooo much, math is the worst subject I participate it. Not grade wise, I have an A. But just doing the questions nearly bore me to death and I just can't understand them without needing someone to help me constantly. Again; thank you. How do I give out a medal?

jimthompson5910 (jim_thompson5910):

just click best response, you're welcome

jimthompson5910 (jim_thompson5910):

maybe try making the questions more exciting so they don't seem so boring?

OpenStudy (anonymous):

I guess, but it's just the numbers itself that bore me out; no matter what kind of information I input into the equation... :(

jimthompson5910 (jim_thompson5910):

i gotcha, well maybe make a game out of it or something? hmm not sure

OpenStudy (anonymous):

There's two more questions I need help with, can you please go over them with me?

OpenStudy (anonymous):

and I guess that makes sense?

jimthompson5910 (jim_thompson5910):

it's just a random thought lol, but sure what's the question

OpenStudy (anonymous):

Ryan’s teacher asked him to graph circle E with equation (x – 5)2 + y2 = 25. Below is the graph he submitted. Circle E that contains points 30, 0; 5, negative 25; negative 20, 0; 5, 25. What, if anything, did Ryan do wrong when graphing this equation? He graphed a circle with a radius of 25 instead of a radius of 5. He graphed a circle with center at (5, 0) instead of at (-5, 0). He graphed a circle with center at (5, 0) instead of at (25, 0). Ryan’s graph is accurate.

jimthompson5910 (jim_thompson5910):

his radius is too big, if you graph the circle, you'll get a radius of 25 not 5

jimthompson5910 (jim_thompson5910):

his radius should be 5

OpenStudy (anonymous):

thank you so much, ok next one?

OpenStudy (anonymous):

how would you graph it exactly?

jimthompson5910 (jim_thompson5910):

either plot the center (5,0) and draw a circle with a radius of 5 around it or draw a circle through those given points (which is harder to do)

OpenStudy (anonymous):

so plot the center the use a compass with the sizing and draw a circle?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

I see! Okay, next one and last one?

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

Circle T has an equation of x2 + (y + 10)2 = 25. Which graph represents circle T?

OpenStudy (anonymous):

not sure if I can post the actual graphs themselves

OpenStudy (anonymous):

any idea @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you can take screenshots and post them

jimthompson5910 (jim_thompson5910):

it's giving me a cookie error...whatever that is

OpenStudy (anonymous):

how do I take a screen shot on a macbook pro?

jimthompson5910 (jim_thompson5910):

hmm one sec

jimthompson5910 (jim_thompson5910):

have a look at this tutorial http://graphicssoft.about.com/od/screencapturemac/ht/macscreenshot.htm